Two sisters shared a $36 5000 inheritance. Teri invested her share at 9%, while Tammy obtained 8.5% interest. After on year, their combined interest was 3200. How much did each sister invest?
Answer by mananth(16946)

I think the inheritance must be $36,500 and not 365,000
Teri 9.00% per annum ---x
Tammy 8.50% per annum ---y

x+y= 36500 ------------------------1
9.00%x+8.50%y= 3200
Multiply by 100
9x+8.5y=320000.00 --------2

Multiply (1) by -9
we get

-9 x -9 y= -328500.00

Add this to (2)

0 x -0.5 y= = -8500

divide by -0.5

y=17000 investment at 8.50%
Balance 19500 investment at 9.00%
CHECK
17000 --------- 8.50% ------- 1445.00
19500 ------- 9.00% ------- 1755.00
Total -------------- 3200.00